OpenStudy (anonymous):

Find the product (2x-1)(x-1)

OpenStudy (anonymous):

yolo solo the hobo

OpenStudy (xapproachesinfinity):

FOIL!

OpenStudy (xapproachesinfinity):

Do you know how to distribute

OpenStudy (anonymous):

no she dont

OpenStudy (xapproachesinfinity):

\((a+b)(c+d)=ac+ad+bc+bd\)

OpenStudy (anonymous):

yea

OpenStudy (xapproachesinfinity):

surely you have seen this many times?

OpenStudy (anonymous):

yes! i just don't know which variable each is

OpenStudy (xapproachesinfinity):

what is \(a*a=?\)

OpenStudy (anonymous):

2a

OpenStudy (anonymous):

1a

OpenStudy (xapproachesinfinity):

incorrect?

OpenStudy (xapproachesinfinity):

that is \(a+a=2a\)

OpenStudy (anonymous):

oh okay??

OpenStudy (xapproachesinfinity):

\(a*a=a^2\) right?

OpenStudy (anonymous):

yes! thats what i meant to put! i forgot how to do the ^

OpenStudy (xapproachesinfinity):

well addition is different than multiplication xD

OpenStudy (anonymous):

lol I know

OpenStudy (xapproachesinfinity):

okay! so what is \(2x*x=?\)

OpenStudy (anonymous):

2^2

OpenStudy (xapproachesinfinity):

first let's apply what we have written above okay \((2x-1)(x-1)=2x*x+2x*(-1)+-(1)*x+(-1)*(-1)\)

OpenStudy (xapproachesinfinity):

I used distributive property that i wrote above with a,b,c,d same thing with our case!

OpenStudy (xapproachesinfinity):

No wrong \(2x*x=2*(x*x)=2x^2\)

OpenStudy (xapproachesinfinity):

remember i wrote \(a*a=a^2\) same with \(x\)

OpenStudy (anonymous):

yolo

OpenStudy (anonymous):

lol I am so confused

OpenStudy (anonymous):

its causee your a blonde

OpenStudy (anonymous):

okkkk

OpenStudy (xapproachesinfinity):

hehe! this is all just like playing a game they give us a definition and property, and we now have something to play with haha the game is \((a+b)(c+d)=ac+ad+bc+bd\) that's rule #1 rule #2 \(a*a=a^2\) multiplying a number by it self rule #3 \((-)*(+)=(-)\\ (+)*(+)=(+)\\ (-)*(-)=(+)\) this * i meant multiplication okay

OpenStudy (xapproachesinfinity):

Now they one us to play this game knowing this facts!

OpenStudy (xapproachesinfinity):

if i tell about a card game rules you would understand the rules first then play the the card game^_^ if i didn't give any rule you will not be able to play it right?

OpenStudy (xapproachesinfinity):

seems my effort and energy was wasted! but anyway... good luck